OpenStudy (anonymous):

Given the function g(x) = −4x + 5, compare and contrast g(2) and g(−4). Choose the statement that is true concerning these two values. The value of g(2) is larger than the value of g(−4). The value of g(2) is smaller than the value of g(−4). The value of g(2) is the same as the value of g(−4). The values of g(2) and g(−4) cannot be compared.

OpenStudy (anonymous):

All you need to do is to plug in 2 and -4.

OpenStudy (cwrw238):

g(2) means the value g(x) when x = 2 so g(2) = -4(2) + 5 = -8 + 5 = -1 now find g(-4) in the same way and you'll be able to work out which statement is true

OpenStudy (cwrw238):

oops sorry -8 + 5 = -3

OpenStudy (cwrw238):

do you follow this?